OpenStudy (anonymous):

solve 0.4(x+0.7)=0.6x(x-42)

OpenStudy (anonymous):

Answer is on the bottom hope it is what you need http://mathway.com/answer.aspx?p=calg?p=0.4(x+0.7)SMB010.6x(x-42)?p=1?p=?p=?p=?p=?p=0?p=?p=0?p=?p=

OpenStudy (anonymous):

I need the steps on how to get the answer

OpenStudy (anonymous):

this website has a little more http://www.wolframalpha.com/input/?i=0.4%28x%2B0.7%29%3D0.6x%28x-42%29

jhonyy9 (jhonyy9):

0,4x+0,28=0,6x2-25,2x multiply both sides with 100 and will get 40x+28=60x2-2520x 60x2-2560x-28=0 simplify by 4 15x2-640x-7=0